OpenStudy (anonymous):

You throw a stone vertically upward with an initial speed of 6.0 m/s from the 3rd story office window. if the window is 12 m from the ground fine the time the stone is in flight and the speed of the stone just before it hits the ground

OpenStudy (popitree):

v = u - gt, so when u throw upwards , at one time v = 0, so u will get t for upward travel, then u can also find the distance, now it is like from tat distance someone just dropped the stone, so calculate how long it will take for the stone to travel (12+x) distance.

OpenStudy (anonymous):

Start by finding out how much time it takes it to reach its highest point (velocity equals 0 there)

OpenStudy (anonymous):

\[v_f=0\] \[0=v_i+a \times t\]

OpenStudy (popitree):

gravity against direction of going[when threw up], so 0 = vi - gt